Using Boolean algebra, simplify the following into the simplest SOP expressions you can. SHOW ALL STEPS.

(A+B)(A'+B)=

A'(A+B)=

(A XOR B)'=

A' + AC=
Answer #1

(1) (A+B)(A'+B) = AA'+AB+A'B+BB

= 0+AB+A'B+B (since AA' = 0, BB = B)

= AB+A'B

= (A+A')B (since A+A' = 1)

(A+B)(A'+B) = B

(2) A'(A+B) = A'A+A'B

= A'B (since AA' = 0)

(3) (A XOR B)' = AB'+A'B

(4) A'+AC = (A'+A)(A'+C) ( since Distributive law x+x'y = (x+x')(x+y) )

= A'+C